### v1

@@
11email: tshenar@tau.ac.il 22institutetext: ESO - European Southern Observatory, Karl-Schwarzschild-Strasse 2, 85748 Garching bei München, Germany 33institutetext: Institute of Astronomy, KU Leuven, Celestijnenlaan 200D, 3001 Leuven, Belgium 44institutetext: Department of Physics & Astronomy, Hounsfield Road, University of Sheffield, Sheffield, S3 7RH, United Kingdom 55institutetext: Instituto de Astrofísica de Canarias, C. Vía Láctea, s/n, 38205 La Laguna, Santa Cruz de Tenerife, Spain 66institutetext: Universidad de La Laguna, Dpto. Astrofísica, Av. Astrofśico Francisco Sánchez, 38206 La Laguna, Santa Cruz de Tenerife, Spain 77institutetext: Escola de Ciências e Tecnologia, Universidade Federal do Rio Grande do Norte, Natal, RN 59072-970, Brazil 88institutetext: Zentrum für Astronomie der Universität Heidelberg, Astronomisches Rechen-Institut, Mönchhofstr. 12-14, 69120 Heidelberg, Germany 99institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Newcastle upon Tyne, NE1 7RU, UK 1010institutetext: Heidelberger Institut für Theoretische Studien, Schloss-Wolfsbrunnenweg 35, 69118 Heidelberg, Germany 1111institutetext: Universität Heidelberg, Department of Physics and Astronomy, Im Neuenheimer Feld 226, 69120 Heidelberg, Germany 1212institutetext: Royal Observatory of Belgium, Avenue Circulaire/Ringlaan 3, B-1180 Brussels, Belgium 1313institutetext: Anton Pannekoek Institute for Astronomy, University of Amsterdam, Science Park 904, 1098 XH Amsterdam, the Netherlands 1414institutetext: Max-Planck-Institute for Astrophysics, Karl-Schwarzschild-Strasse 1, 85748 Garching, Germany 1515institutetext: European Space Agency (ESA), ESA Office, Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 1616institutetext: ICREA, Pg. Lluís Companys 23, E08010 Barcelona, Spain 1717institutetext: Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (IEEC-UB), Martí Franquès 1, E08028 Barcelona, Spain 1818institutetext: Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, United Kingdom 1919institutetext: Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 2020institutetext: Institute of Science and Technology Austria (ISTA), Am Campus 1, 3400 Klosterneuburg, Austria 2121institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2222institutetext: Department of Astronomy and Physics, Saint Mary’s University, 923 Robie Street, Halifax, B3H 3C3, Canada 2323institutetext: Gemini Observatory/NSF’s NOIRLab, Casilla 603, La Serena, Chile 2424institutetext: Center for Computational Astrophysics, Division of Science, National Astronomical Observatory of Japan, 2-21-1, Osawa, Mitaka, Tokyo 181-8588, Japan 2525institutetext: Astronomical Institute, Academy of Sciences of the Czech Republic, Fričova 298, CZ-251 65 Ondřejov, Czech Republic 2626institutetext: Department of Astronomy and Astrophysics, University of Toronto, 50 St. George Street, Toronto, Ontario, M5S 3H4, Canada 2727institutetext: Dublin Institute for Advanced Studies, DIAS Dunsink Observatory, Dunsink Lane, Dublin 15, Ireland 2828institutetext: Centro de Astrobiología (CSIC-INTA), campus ESAC, camino bajo del castillo s/n, 28 692 Villanueva de la Cañada, Spain 2929institutetext: School of Physics and Astronomy, Monash University, Clayton VIC 3800, Australia 3030institutetext: ARC Centre of Excellence for Gravitational-wave Discovery (OzGrav), Melbourne, Australia 3131institutetext: IAASARS, National Observatory of Athens, GR-15236, Penteli, Greece 3232institutetext: Institute of Astrophysics, FORTH, GR-71110, Heraklion, Greece 3333institutetext: Centro de Astrobiología (CSIC-INTA), Ctra. Torrejón a Ajalvir km 4, 28850 Torrejón de Ardoz, Spain 3434institutetext: Institut für Physik und Astronomie, Universität Potsdam, Karl-Liebknecht-Str. 24/25, 14476 Potsdam, Germany 3535institutetext: Lund Observatory, Division of Astrophysics, Department of Physics, Lund University, Box 43, SE-221 00, Lund, Sweden 3636institutetext: Department of Astronomy & Steward Observatory, 933 N. Cherry Ave., Tucson, AZ 85721, USA 3737institutetext: Observatório Nacional, R. Gen. José Cristino, 77 - Vasco da Gama, Rio de Janeiro - RJ, 20921-400, Brazil 3838institutetext: Institute of Astronomy, Faculty of Physics, Astronomy and Informatics, Nicolaus Copernicus University, Grudziadzka 5, 87-100 Torun, Poland 3939institutetext: Department of Particle Physics and Astrophysics, Weizmann Institute of Science, Rehovot 7610001, Israel 4040institutetext: Lennard-Jones Laboratories, Keele University, ST5 5BG, UK 4141institutetext: Center for Computational Astrophysics, Flatiron Institute, New York, NY 10010, USA 4242institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany 4343institutetext: Armagh Observatory, College Hill, Armagh, BT61 9DG, Northern Ireland, UK
@@
Binarity at LOw Metallicity (BLOeM)††thanks: Based on observations collected at the European Southern Observatory under ESO program ID 112.25W2.,††thanks: Tables … are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@
Surveys in the Milky Way and Large Magellanic Cloud revealed that the majority of massive stars will interact with companions during their lives. However, knowledge of the binary properties of massive stars at low metallicity, which approaches the conditions of the Early Universe, remains sparse. We present the Binarity at LOw Metallicity (BLOeM) campaign – an ESO large programme designed to obtain 25 epochs of spectroscopy for 929 massive stars in the Small Magellanic Cloud (SMC) – the lowest metallicity conditions in which multiplicity is probed to date (). BLOeM will provide (i) the binary fraction, (ii) the orbital configurations of systems with periods yr, (iii) dormant black-hole binary candidates (OB+BH), and (iv) a legacy database of physical parameters of massive stars at low metallicity.
@@
The stars are observed with the LR02 setup of the giraffe instrument of the Very Large Telescope (3960 – 4570Å resolving power ; typical signal-to-noise ratio S/N) . This paper utilises the first 9 epochs obtained over a three-month time. We describe the survey and data reduction, perform a spectral classification of the stacked spectra, and construct a Hertzsprung-Russell diagram of the sample via spectral-type and photometric calibrations.
@@
The sample covers spectral types from O4 to F5, spanning the effective temperature and luminosity ranges and and initial masses . It comprises 159 O-type stars, 324 early B-type (B0–3) dwarfs and giants (luminosity classes V–III), 309 early B-type supergiants (II–I), and 137 late-type supergiants. At least 75 stars are Oe/Be stars: 20 O-type and 55 B-type (13% and 10% of the respective samples). In addition, it includes four high-mass X-ray binaries, three stars resembling luminous blue variables, two bloated stripped-star candidates, two candidate magnetic stars, and 74 eclipsing binaries.
@@
Massive stars (), typically classified as OB-type stars on the main sequence, play an increasingly important role in modern astrophysics, with direct links to stellar dynamics, stellar transients, and gravitational-wave (GW) astrophysics (Portegies Zwart & Verbunt, 1996; Langer, 2012; Woosley, 2017; Wang et al., 2020; Marchant & Bodensteiner, 2023). Spectroscopic and interferometric surveys in the Milky Way (MW) and the Large Magellanic Cloud (LMC) have shown that stellar multiplicity is common among massive stars (e.g., Abt, 1983; Kobulnicky & Fryer, 2007; Mason et al., 2009; Sana et al., 2012, 2014; Bordier et al., 2023); over 50% of massive stars are expected to interact with a companion during their lifetime (Sana et al., 2012, 2013; Stegmann et al., 2022; Offner et al., 2023; Kummer et al., 2023). Such interactions dramatically alter the evolutionary paths and final fates of massive stars, resulting in a plethora of astronomical phenomena such as stripped helium stars and Wolf-Rayet (WR) stars (Paczyński, 1967; Shenar et al., 2016; Götberg et al., 2018; Pauli et al., 2022a; Drout et al., 2023), rapidly rotating stars with decretion disks (Oe/Be stars; Pols et al., 1991; Rivinius et al., 2013; de Mink et al., 2013; Wang et al., 2017; Bodensteiner et al., 2020; Britavskiy et al., 2023; Renzo & Götberg, 2021), stellar mergers and magnetic stars (Ferrario et al., 2009; de Mink et al., 2014; Schneider et al., 2019; Shenar et al., 2023; Frost et al., 2024), single-degenerate binaries and high-mass X-ray binaries (Corral-Santana et al., 2016; Shenar et al., 2022a; Mahy et al., 2022), and GW sources (de Mink & Mandel, 2016; Marchant et al., 2016; Tauris et al., 2017; Mandel & Broekgaarden, 2022; Mandel & Farmer, 2022).
@@
Of special interest is massive-star research in low metallicity (low ) environments, which reflect the conditions prevalent in the distant Universe. An increasing number of transients, such as long-duration -ray bursts (LGRBs; Yoon & Langer, 2005; Woosley & Heger, 2006), superluminous supernovae (SNe; Quimby et al., 2011; Gal-Yam, 2012), broad-lined type Ic supernovae (Modjaz et al., 2008), and pair-instability SNe (Barkat et al., 1967; Fryer et al., 2001; Langer et al., 2007; Woosley, 2017; Farmer et al., 2019) are thought to be associated mainly or exclusively with low- conditions. Similarly, the bulk of black-hole (BH) mergers observed with the LIGO-Virgo-KAGRA collaboration are thought to originate in low metallicity conditions (Abbott et al., 2016; Giacobbo et al., 2018; Klencki et al., 2018).
@@
Modern investigations expose deficiencies in our understanding of massive stars at low . For example, the rates and mass distribution of BH mergers defy original expectations (Broekgaarden et al., 2021; Mandel & Broekgaarden, 2022; van Son et al., 2022), and observables such as the rate of LGRBs (Graham & Fruchter, 2017; Chen et al., 2017) and the fraction of Oe/Be stars and Be X-ray binaries (Haberl & Sturm, 2016; Schootemeijer et al., 2022) as a function of are not reproduced by contemporary models (Graham & Fruchter, 2017; Chen et al., 2017; Hastings et al., 2021). Such discrepancies are likely related to insufficient knowledge of massive-star evolution at low , or to a false extrapolation of the initial conditions of massive stars (e.g., binary fraction, orbital configurations) to low . The recent detection of a BH with a low-metallicity companion through Gaia astrometry provides additional evidence suggesting that low- environments are crucial for the formation of massive BHs (Gaia Collaboration et al., 2024).
@@
While it has been shown that for solar-type stars binary properties can depend on the natal metallicity, this remains only a prediction for massive stars (Kroupa, 2001; Saigo et al., 2004; Machida, 2008; Marks et al., 2012; Moe et al., 2019; Price-Whelan et al., 2020). To mitigate this, spectroscopic monitoring surveys of massive-star populations at different environments are needed, which are sensitive to the regime of binary interactions (i.e. orbital periods ). A notable spectroscopic campaign in this context was the VLT-FLAMES Tarantula Survey (VFTS; PI: Evans), which monitored massive stars in the Tarantula nebula of the LMC (Evans et al., 2011), with a metallicity of . The VFTS survey yielded a comparable intrinsic binary fraction for OB-type stars in the LMC (50-60%; Sana et al. 2013; Dunstall et al. 2015) to that observed in different Galactic environments (50-70%, e.g., Sana et al., 2012; Kiminki & Kobulnicky, 2012; Banyard et al., 2022). The follow-up Tarantula Massive Binary Monitoring (TMBM; PI: Sana) and B-type Binary Characterisation (BBC; PI: Taylor) programmes also revealed an overall similar distribution of orbital parameters and mass ratios to Galactic samples (Almeida et al., 2017; Villaseñor et al., 2021; Shenar et al., 2022b; Mahy et al., 2020b). However, the LMC metallicity only differs by a factor of from that of the MW.
@@
The Small Magellanic Cloud (SMC) is a neighbouring dwarf galaxy with (Hunter et al., 2007) located about 62 kpc from Earth (Graczyk et al., 2020), hosting thousands of massive stars (Humphreys & McElroy, 1984). It had a star formation peak Myr ago (Antoniou et al., 2010; Rubele et al., 2015; Schootemeijer et al., 2021), potentially triggered by a collision with the LMC about 100-150 Myr ago (e.g. Zivick et al., 2018). While galaxies of lower metal content exist in the Local Group (e.g., Sextans A, Skillman et al. 1989; Lorenzo et al. 2022; Leo P, McQuinn et al. 2015; Evans et al. 2019; Telford et al. 2023), the SMC is the only galaxy in which a large sample of massive stars at low can currently be resolved and spectroscopically monitored with sufficient spectral resolution and signal-to-noise (S/N). Previous or ongoing surveys addressed aspects related to the SMC massive-star contents (Humphreys & McElroy, 1984; Evans et al., 2006; Martayan et al., 2007; Evans et al., 2004; Schootemeijer et al., 2021), stellar winds and mass-loss (Ramachandran et al., 2019; Vink et al., 2023) and runaway status (RIOTS4, Lamb et al., 2016), but multiplicity has been largely neglected beyond analyses of selected eclipsing binaries (Hilditch et al., 2005), clusters (Dufton et al., 2019; Bodensteiner et al., 2021), and individual objects of interest (e.g., Pauli et al., 2022b). Moe & Di Stefano (2013) investigated the frequency of eclipsing massive binaries among B-type stars in the SMC, LMC, and MW, and found no significant differences between the populations. However, their study was limited to the period range d, and is generally bias-dominated given the small fraction of eclipsing binaries. Further information regarding multiplicity in the SMC is not known.
@@
The need to establish the multiplicity of massive stars at low is not the only reason to monitor massive stars in the SMC. Recent spectroscopic monitoring of massive stars in the MW (Mahy et al., 2022) and the LMC (Shenar et al., 2022b, a) uncovered the tip of the iceberg of a new population of massive single-degenerate binaries: X-ray dormant OB+BH binaries (e.g., Giesers et al., 2018). Such binaries yield precious constraints on core-collapse mechanisms and the presence of SN explosions and possible natal kicks during the collapse into BHs (Mirabel & Rodrigues, 2003; Renzo et al., 2019; Atri et al., 2019; Langer et al., 2020; Banagiri et al., 2023). For example, Vigna-Gómez et al. (2024) recently used the dormant OB+BH binary VFTS 243 to derive a natal kick of 4 km and an ejection of 0.3 neutrino mass during the collapse of the progenitor. The Gaia mission will likely uncover dozens more OB+BH binaries in the MW via high-precision astrometry (Mashian & Loeb, 2017; Breivik et al., 2017; Janssens et al., 2022, 2023), though so far only low-mass stars with BH companions were discovered (El-Badry et al., 2023b, a; Shahaf et al., 2023; Gaia Collaboration et al., 2024). The formation scenarios of these objects are still debated, and it is possible that they originated via dynamical captures, making them less constraining for SN physics (Rastello et al., 2023; El-Badry, 2024; Marín Pina et al., 2024). In any case, Gaia will not uncover extragalactic OB+BH binaries. Finding the first dormant OB+BH binaries in the SMC via spectroscopic monitoring of massive stars has the potential to yield unprecedented constraints on BH formation at low .
@@
Motivated by these reasons, we initiated a novel spectroscopic monitoring survey of a large population of massive stars in the SMC. The Binarity at LOw Metallicity (BLOeM) campaign is a European Southern Observatory (ESO) Large Programme (PI: Shenar, dPI: Bodensteiner; ID: 112.25R7) scheduled for 2023 – 2025. Relying on 116 hr of observing time with the Fibre Large Array Multi Element Spectrograph (FLAMES; Pasquini et al. 2002) of the Very Large Telescope (VLT), the survey is underway and assembling 25 epochs of spectroscopy for 929 massive stars for a total baseline of two years (four semesters). The survey will enable a full characterisation of the binary fraction and orbital parameters of the stars up to orbital periods of a few years and mass ratios as low as , discovery of dormant OB+BH binaries, and a complete analysis of the binary and single-star content of the sample.

### v2

@@
11email: tshenar@tau.ac.il 22institutetext: ESO - European Southern Observatory, Karl-Schwarzschild-Strasse 2, 85748 Garching bei München, Germany 33institutetext: Institute of Astronomy, KU Leuven, Celestijnenlaan 200D, 3001 Leuven, Belgium 44institutetext: Department of Physics & Astronomy, Hounsfield Road, University of Sheffield, Sheffield, S3 7RH, United Kingdom 55institutetext: Instituto de Astrofísica de Canarias, C. Vía Láctea, s/n, 38205 La Laguna, Santa Cruz de Tenerife, Spain 66institutetext: Universidad de La Laguna, Dpto. Astrofísica, Av. Astrofśico Francisco Sánchez, 38206 La Laguna, Santa Cruz de Tenerife, Spain 77institutetext: Escola de Ciências e Tecnologia, Universidade Federal do Rio Grande do Norte, Natal, RN 59072-970, Brazil 88institutetext: Zentrum für Astronomie der Universität Heidelberg, Astronomisches Rechen-Institut, Mönchhofstr. 12-14, 69120 Heidelberg, Germany 99institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Newcastle upon Tyne, NE1 7RU, UK 1010institutetext: Heidelberger Institut für Theoretische Studien, Schloss-Wolfsbrunnenweg 35, 69118 Heidelberg, Germany 1111institutetext: Universität Heidelberg, Department of Physics and Astronomy, Im Neuenheimer Feld 226, 69120 Heidelberg, Germany 1212institutetext: Royal Observatory of Belgium, Avenue Circulaire/Ringlaan 3, B-1180 Brussels, Belgium 1313institutetext: Anton Pannekoek Institute for Astronomy, University of Amsterdam, Science Park 904, 1098 XH Amsterdam, the Netherlands 1414institutetext: Max-Planck-Institute for Astrophysics, Karl-Schwarzschild-Strasse 1, 85748 Garching, Germany 1515institutetext: European Space Agency (ESA), ESA Office, Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 1616institutetext: ICREA, Pg. Lluís Companys 23, E08010 Barcelona, Spain 1717institutetext: Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (IEEC-UB), Martí Franquès 1, E08028 Barcelona, Spain 1818institutetext: Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, United Kingdom 1919institutetext: Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 2020institutetext: Institute of Science and Technology Austria (ISTA), Am Campus 1, 3400 Klosterneuburg, Austria 2121institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2222institutetext: Department of Astronomy and Physics, Saint Mary’s University, 923 Robie Street, Halifax, B3H 3C3, Canada 2323institutetext: Gemini Observatory/NSF’s NOIRLab, Casilla 603, La Serena, Chile 2424institutetext: Center for Computational Astrophysics, Division of Science, National Astronomical Observatory of Japan, 2-21-1, Osawa, Mitaka, Tokyo 181-8588, Japan 2525institutetext: Astronomical Institute, Academy of Sciences of the Czech Republic, Fričova 298, CZ-251 65 Ondřejov, Czech Republic 2626institutetext: Department of Astronomy and Astrophysics, University of Toronto, 50 St. George Street, Toronto, Ontario, M5S 3H4, Canada 2727institutetext: Dublin Institute for Advanced Studies, DIAS Dunsink Observatory, Dunsink Lane, Dublin 15, Ireland 2828institutetext: Centro de Astrobiología (CSIC-INTA), campus ESAC, camino bajo del castillo s/n, 28 692 Villanueva de la Cañada, Spain 2929institutetext: School of Physics and Astronomy, Monash University, Clayton VIC 3800, Australia 3030institutetext: ARC Centre of Excellence for Gravitational-wave Discovery (OzGrav), Melbourne, Australia 3131institutetext: IAASARS, National Observatory of Athens, GR-15236, Penteli, Greece 3232institutetext: Institute of Astrophysics, FORTH, GR-71110, Heraklion, Greece 3333institutetext: Centro de Astrobiología (CSIC-INTA), Ctra. Torrejón a Ajalvir km 4, 28850 Torrejón de Ardoz, Spain 3434institutetext: Institut für Physik und Astronomie, Universität Potsdam, Karl-Liebknecht-Str. 24/25, 14476 Potsdam, Germany 3535institutetext: McWilliams Center for Cosmology & Astrophysics, Department of Physics, Carnegie Mellon University, Pittsburgh, PA 15213, USA 3636institutetext: Lund Observatory, Division of Astrophysics, Department of Physics, Lund University, Box 43, SE-221 00, Lund, Sweden 3737institutetext: Department of Astronomy & Steward Observatory, 933 N. Cherry Ave., Tucson, AZ 85721, USA 3838institutetext: Observatório Nacional, R. Gen. José Cristino, 77 - Vasco da Gama, Rio de Janeiro - RJ, 20921-400, Brazil 3939institutetext: Institute of Astronomy, Faculty of Physics, Astronomy and Informatics, Nicolaus Copernicus University, Grudziadzka 5, 87-100 Torun, Poland 4040institutetext: Department of Particle Physics and Astrophysics, Weizmann Institute of Science, Rehovot 7610001, Israel 4141institutetext: Lennard-Jones Laboratories, Keele University, ST5 5BG, UK 4242institutetext: Center for Computational Astrophysics, Flatiron Institute, New York, NY 10010, USA 4343institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany 4444institutetext: Armagh Observatory, College Hill, Armagh, BT61 9DG, Northern Ireland, UK
@@
Binarity at LOw Metallicity (BLOeM)††thanks: Based on observations collected at the European Southern Observatory under ESO program ID 112.25W2.,††thanks: Table LABEL:tab:SpT is available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@
Surveys in the Milky Way and Large Magellanic Cloud have revealed that the majority of massive stars will interact with companions during their lives. However, knowledge of the binary properties of massive stars at low metallicity, and therefore in conditions approaching those of the Early Universe, remain sparse. We present the Binarity at LOw Metallicity (BLOeM) campaign, an ESO large programme designed to obtain 25 epochs of spectroscopy for 929 massive stars in the Small Magellanic Cloud, allowing us to probe multiplicity in the lowest-metallicity conditions to date (). BLOeM will provide (i) the binary fraction, (ii) the orbital configurations of systems with periods of yr, (iii) dormant black-hole binary candidates (OB+BH), and (iv) a legacy database of physical parameters of massive stars at low metallicity.
@@
Main sequence (OB-type) and evolved (OBAF-type) massive stars are observed with the LR02 setup of the giraffe instrument of the Very Large Telescope (3960 – 4570Å resolving power ; typical signal-to-noise ratio(S/N) ) . This paper utilises the first nine epochs obtained over a three-month time period. We describe the survey and data reduction, perform a spectral classification of the stacked spectra, and construct a Hertzsprung-Russell diagram of the sample via spectral-type and photometric calibrations.
@@
Our detailed classification reveals that the sample covers spectral types from O4 to F5, spanning the effective temperature and luminosity ranges and and initial masses of . The sample comprises 159 O-type stars, 331 early B-type (B0–3) dwarfs and giants (luminosity classes V–III), 303 early B-type supergiants (II–I), and 136 late-type BAF supergiants. At least 82 stars are OBe stars: 20 O-type and 62 B-type (13% and 11% of the respective samples). In addition, the sample includes 4 high-mass X-ray binaries, 3 stars resembling luminous blue variables, 2 bloated stripped-star candidates, 2 candidate magnetic stars, and 74 eclipsing binaries.
@@
Massive stars (), typically classified as OB-type stars on the main sequence, play an increasingly important role in modern astrophysics, with direct links to stellar dynamics, stellar transients, and gravitational-wave (GW) astrophysics (Portegies Zwart & Verbunt, 1996; Langer, 2012; Woosley, 2017; Wang et al., 2020; Marchant & Bodensteiner, 2023). Spectroscopic and interferometric surveys in the Milky Way (MW) and the Large Magellanic Cloud (LMC) have shown that stellar multiplicity is common among massive stars (e.g. Abt, 1983; Kobulnicky & Fryer, 2007; Mason et al., 2009; Sana et al., 2012, 2014; Bordier et al., 2024); over 50% of massive stars are expected to interact with a companion during their lifetime (Sana et al., 2012, 2013; Stegmann et al., 2022; Offner et al., 2023; Kummer et al., 2023). Such interactions dramatically alter the evolutionary paths and final fates of massive stars, resulting in a plethora of astronomical phenomena such as stripped helium stars and Wolf-Rayet (WR) stars (Paczyński, 1967; Shenar et al., 2016; Götberg et al., 2018; Pauli et al., 2022a; Drout et al., 2023), rapidly rotating stars with decretion disks (OBe stars; Pols et al., 1991; Rivinius et al., 2013; de Mink et al., 2013; Wang et al., 2017; Bodensteiner et al., 2020; Britavskiy et al., 2023; Renzo & Götberg, 2021), stellar mergers and magnetic stars (Ferrario et al., 2009; de Mink et al., 2014; Schneider et al., 2019; Shenar et al., 2023; Frost et al., 2024), single-degenerate binaries and high-mass X-ray binaries (Corral-Santana et al., 2016; Shenar et al., 2022a; Mahy et al., 2022), and GW sources (de Mink & Mandel, 2016; Marchant et al., 2016; Tauris et al., 2017; Mandel & Broekgaarden, 2022; Mandel & Farmer, 2022).
@@
Of special interest is massive-star research in low-metallicity (low ) environments, which reflect the conditions prevalent in the distant Universe. An increasing number of transients, such as long-duration -ray bursts (LGRBs; Yoon & Langer, 2005; Woosley & Heger, 2006), superluminous supernovae (SNe; Quimby et al., 2011; Gal-Yam, 2012), broad-lined type Ic supernovae (Modjaz et al., 2008), and pair-instability SNe (Barkat et al., 1967; Fryer et al., 2001; Langer et al., 2007; Woosley, 2017; Farmer et al., 2019), are thought to be associated mainly or exclusively with low- conditions. Similarly, the bulk of black-hole (BH) mergers observed with the LIGO-Virgo-KAGRA collaboration are thought to originate in low-metallicity conditions (Abbott et al., 2016; Giacobbo et al., 2018; Klencki et al., 2018).
@@
Modern investigations expose deficiencies in our understanding of massive stars at low . For example, the rates and mass distribution of BH mergers defy original expectations (Broekgaarden et al., 2021; Mandel & Broekgaarden, 2022; van Son et al., 2022), and observables such as the rate of LGRBs (Graham & Fruchter, 2017; Chen et al., 2017) and the fraction of OBe stars and Be X-ray binaries (Haberl & Sturm, 2016; Schootemeijer et al., 2022) as a function of are not reproduced by contemporary models (Graham & Fruchter, 2017; Chen et al., 2017; Hastings et al., 2021). Such discrepancies are likely related to insufficient knowledge of massive-star evolution at low , or to a false extrapolation of the initial conditions of massive stars (e.g. binary fraction, orbital configurations) to low . The recent detection of a BH with a low-metallicity companion through Gaia astrometry provides additional evidence suggesting that low- environments are crucial for the formation of massive BHs (Gaia Collaboration et al., 2024).
@@
While it has been shown that the binary properties of solar-type stars can depend on natal metallicity, this remains a prediction for massive stars (Kroupa, 2001; Saigo et al., 2004; Machida, 2008; Marks et al., 2012; Moe et al., 2019; Price-Whelan et al., 2020). To mitigate this, we need spectroscopic monitoring surveys of massive-star populations at different environments sensitive to the regime of binary interactions (i.e. orbital periods ). A notable spectroscopic campaign in this context was the VLT-FLAMES Tarantula Survey (VFTS; PI: Evans), which monitored about 1000 massive stars in the Tarantula nebula of the LMC (Evans et al., 2011), which has a metallicity of . The VFTS survey yielded a comparable intrinsic binary fraction for OB-type stars in the LMC (50-60%; Sana et al. 2013; Dunstall et al. 2015) to that observed in different Galactic environments (50-70%, e.g. Sana et al., 2012; Kiminki & Kobulnicky, 2012; Banyard et al., 2022; Guo et al., 2022). The follow-up Tarantula Massive Binary Monitoring (TMBM; PI: Sana) and B-type Binary Characterisation (BBC; PI: Taylor) programmes also revealed an overall similar distribution of orbital parameters and mass ratios to Galactic samples (Almeida et al., 2017; Villaseñor et al., 2021; Shenar et al., 2022b; Mahy et al., 2020b). However, the LMC metallicity only differs by a factor of from that of the MW.
@@
The Small Magellanic Cloud (SMC) is a neighbouring dwarf galaxy with (Hunter et al., 2007) located about 62 kpc from Earth (Graczyk et al., 2020), hosting thousands of massive stars (Humphreys & McElroy, 1984). It had a star-formation peak Myr ago (Antoniou et al., 2010; Rubele et al., 2015; Schootemeijer et al., 2021), potentially triggered by a collision with the LMC about 100-150 Myr ago (e.g. Zivick et al., 2018). While galaxies of lower metal content exist in the Local Group (e.g. Sextans A, Skillman et al. 1989; Lorenzo et al. 2022; Leo P, McQuinn et al. 2015; Evans et al. 2019; Telford et al. 2023), the SMC is the only galaxy in which a large sample of massive stars at low can currently be resolved and spectroscopically monitored with sufficient spectral resolution and signal-to-noise ratio (S/N). Previous or ongoing surveys addressed aspects related to the SMC massive-star contents (Humphreys & McElroy, 1984; Evans et al., 2006; Martayan et al., 2007; Evans et al., 2004; Schootemeijer et al., 2021), stellar winds and mass-loss (Ramachandran et al., 2019; Vink et al., 2023), and runaway status (RIOTS4, Lamb et al., 2016), but multiplicity has been largely neglected beyond analyses of selected eclipsing binaries (Hilditch et al., 2005), clusters (Dufton et al., 2019; Bodensteiner et al., 2021), and individual objects of interest (e.g. Pauli et al., 2022b). Moe & Di Stefano (2013) investigated the frequency of eclipsing massive binaries among B-type stars in the SMC, LMC, and MW, and found no significant differences between the populations. However, their study was limited to the period range d, and is generally bias-dominated given the small fraction of eclipsing binaries. There is no further information available regarding multiplicity in the SMC.
@@
The need to establish the multiplicity of massive stars at low is not the only reason to monitor massive stars in the SMC. Recent spectroscopic monitoring of massive stars in the MW (Mahy et al., 2022) and the LMC (Shenar et al., 2022b, a) uncovered the ‘tip of the iceberg’ of a new population of massive single-degenerate binaries: X-ray dormant OB+BH binaries (e.g. Giesers et al., 2018). Such binaries yield precious constraints on core-collapse mechanisms and the presence of SN explosions and possible natal kicks during the collapse into BHs (Mirabel & Rodrigues, 2003; Renzo et al., 2019; Atri et al., 2019; Langer et al., 2020; Banagiri et al., 2023). For example, Vigna-Gómez et al. (2024) recently used the dormant OB+BH binary VFTS 243 to derive a natal kick of 4 km and an ejection of 0.3 neutrino mass during the collapse of the progenitor. The Gaia mission will likely uncover dozens more OB+BH binaries in the MW via high-precision astrometry (Mashian & Loeb, 2017; Breivik et al., 2017; Janssens et al., 2022, 2023), though so far only low-mass stars with BH companions have been discovered (El-Badry et al., 2023b; Chakrabarti et al., 2023; El-Badry et al., 2023a; Shahaf et al., 2023; Gaia Collaboration et al., 2024). The formation scenarios of these objects are still debated, and it is possible that they originate from dynamical captures, making them less useful for constraints on SN physics (Rastello et al., 2023; El-Badry, 2024; Marín Pina et al., 2024). In any case, Gaia will not uncover extragalactic OB+BH binaries. Finding the first dormant OB+BH binaries in the SMC via spectroscopic monitoring of massive stars has the potential to yield unprecedented constraints on BH formation at low .
@@
Motivated by these objectives, we initiated a novel spectroscopic monitoring survey of a large population of massive stars in the SMC. The Binarity at LOw Metallicity (BLOeM) campaign is a European Southern Observatory (ESO) Large Programme (PI: Shenar, dPI: Bodensteiner; ID: 112.25R7) scheduled for 2023 – 2025. Relying on 116 hr of observing time with the Fibre Large Array Multi Element Spectrograph (FLAMES; Pasquini et al. 2002) of the Very Large Telescope (VLT), the survey is underway and assembling 25 epochs of spectroscopy for 929 massive stars for a total baseline of two years (four semesters). The survey will enable a full characterisation of the binary fraction and the orbital parameters of stars with orbital periods of up to a few years and with mass ratios of as low as ; the discovery of dormant OB+BH binaries; and a complete analysis of the binary and single-star content of the sample.